What would the acceleration of a freely falling body be measured as in an elevatorthat was moving upward with an acceleration of 2 m/s2? Same question if the elevator was moving downward with an acceleration of 2
Answer #1

for upward moving elevator

g' = g + a

g' = 9.8 + 2 = 11.8 m/s^2

======

for downward moving elevator

g' = g - a

g' = 9.8 - 2

g' = 7.8 m/s^2

======

Comment before rate in case any doubt, will reply for sure.. goodluck